two identical spheres carrying charges 9 micro coulomb and 5 micro coulomb respectively are kept in contact and then seperated from each other . how many extra electrons wouyld be present 

Grade:12th pass

1 Answers

Arun
4 years ago
Dear student
 
q = 5-9/2 = – 2 micro coulomb
 
ne = -2 * 10^-6 
 
n = – 2 * 10^-6 /1.6 * 10^-19 = 1.25 * 10^13 electrons